# Adaptive Partition Factor Analysis

### v1

@@ 2.2 Model properties
Identifiability of factor models is an overarching issue. On the one hand, in fact, factor models are appealing in providing possible interpretations and insights on how a small set of latent factors is connected to a multivariate observation with a possible intricate form of dependency. The other side of the medal, however, is that mathematically a factor analytic model is not identifiable in many respect. See [7] [5] [6] for a comprehensive review. For instance, loadings are not identifiable due to rotation, sign and permutation invariance. Rotational invariance refers to transformation of loadings and factors by a orthogonal matrix, sign switching arises because we can simultaneously change the signs of and certain elements without affecting the overall structure while permutation invariance occurs because there is no intrinsic ordering in the factor loading matrix columns.
@@
Identification issues are exacerbated in the MSFA framework of [8] where it may be difficult to disentangle the signal between shared and study-specific factors, a problem known as information switching. For instance, [3] show how a linear combination of study-specific factors, can instead be represented and reincorporated as part of the shared component. Building on these concepts, we provide conditions to identify the shared from the specific factors in our setup. More formally, we give the following definition of information switching.
Definition 2.1.
@@
Let $n$ be the number of units, and $S_{n}$ the number of distinct groups in model (9), i.e. $S_{n}=\left|\cup_{i=1}^{n}\Omega_{i}\right|$. Denote with $\Psi$ the $n\times k$ matrix that stacks in distinct rows all local scale parameters $\psi_{i}=(\psi_{i1},\ldots,\psi_{ik})$ and with $\Psi_{h}$ its generic column with $\Psi_{h}\neq 1_{n}$ for all $h\in\{1,\ldots,k\}$. Define $W_{s}=\Omega_{s}-\Lambda\Lambda^{\top}-\Sigma=\Gamma\text{diag}\{\psi_{s}\}\Gamma^{\top}$, for $s\in\{1,\ldots,S_{n}\}$. The model suffers from information switching if there exist $\tilde{\Gamma}\neq\Gamma$ and $\tilde{\Psi}\neq\Psi$ such that $W_{s}=\tilde{\Gamma}{\text{diag}\{\tilde{\psi}_{s}\}}\tilde{\Gamma}$ for all $s$, with $\tilde{\Psi}_{h}=1_{n}$ for at least one $h$.
@@
It is clear that if there is $\tilde{\Psi}_{h}=1_{n}$, the $h$-th factor could no longer be interpreted as specific and should be moved to the shared part and, consistently, $\Lambda\Lambda^{\top}$ would represent only a fraction of the shared variance. Note that in Definition 2.1, $S_{n}$ does not refer to the number of distinct labels or studies $S$ fixed a priori and embedded in the categorical variables $x_{i}$, but rather it is equal to the number of distinct configurations of the rows of the matrix $\Psi$. The following Theorem ensures the identification in terms of information switching.
Theorem 2.2.
@@
For the model defined in (8), if $\Gamma$ is of full column rank $k<p(p+1)/2$ and $\Gamma\Gamma^{\top}$ is not sparse, then the model is resistant to information switching.
@@
Since each $\psi_{s}$ is a vector of zeroes and ones, we can see that $W_{s}=\Gamma\text{diag}\{\psi_{s}\}\Gamma^{\top}=\Gamma\text{diag}\{\psi_{s}\}\text{diag}\{\psi_{s}\}\Gamma^{\top}.$ Define $\Gamma_{s}=\Gamma\text{diag}\{\psi_{s}\}$. As in any factor analytic decompositions, the elements of $W_{s}$ are identifiable up to multiplication by group-specific rotation matrices $P_{s}$, such that $P_{s}P_{s}^{\top}=I_{k}$ and $|P_{s}|=1$. Hence we can write $W_{s}=\Gamma_{s}P_{s}P_{s}^{\top}\Gamma_{s}^{\top}=\Gamma\text{diag}\{\psi_{s}\}P_{s}P_{s}^{\top}\text{diag}\{\psi_{s}\}\Gamma^{\top}.$ Hence each element $w^{jl}_{s}$ of $W_{s}$, can be written as
$w^{jl}_{s}=\sum_{h=1}^{k}\gamma_{jh}\gamma_{lh}{\bar{\psi}_{hs}\bar{\psi}_{hs}^{\top}},$
where $\bar{\psi}_{hs}$ is the $h$-th row of the product $\text{diag}\{\psi_{s}\}P_{s}$. Note that the product $\bar{\psi}_{hs}\bar{\psi}_{hs}^{\top}$ is either 0 or 1 by construction. Thus we know that there exists $\tilde{\psi}_{hs}=\bar{\psi}_{hs}\bar{\psi}_{hs}^{\top}$ for $h=1,\dots k$ that lead to the same $W_{s}$ for each $s=1,\dots,S_{n}$. Hence, it is sufficient to check whether these $\tilde{\psi}_{s}$ are different from the original $\psi_{s}$ for each $s$. Since we have $S_{n}p(p+1)/2$ equations of the form
$\sum_{h=1}^{k}\gamma_{jh}\gamma_{hl}\tilde{\psi}_{hs}-w^{jl}_{s}=0,$
with $S_{n}k$ unknowns and $k<p(p+1)/2$ by hypothesis, there will be at most one solution. ∎
Based on the identifiability finding of Theorem 2.2, the practical condition that ensures resistance to information switching in APAFA are discussed in Lemma 2.3 and 2.4.
Lemma 2.3.
@@
A key advantage of our model is that the requirement of $k<p(p+1)/2$ is minimal as it is customary in factor analysis to have latent factors of much lower dimension than the observed data ($k<p$). Notably, [3], in different settings, necessitate the dimension of the specific part to be smaller than that of the shared part implicitly considering the study-specific part as a noise or disturbance parameter. Although we are in different settings, where the number and composition of the groups is not pre-specified, we do not restrict $k$ to be also smaller than $d$. This is crucial in many contexts. For instance, in cancer studies, it is possible to study very different types of cancer simultaneously, identifying a range of specific factors that describe the specific characteristics of each type of cancer, while also uncovering a few common factors that are shared across cancers, regardless of the organ or tissue affected possibly pertaining to fundamental mechanisms of oncogenesis.
Notably, the specific structure induced by sparsity in the $\psi_{s}$ not only ensures identifiability in terms of information switching but also solves the usual rotational ambiguity, under certain conditions. This is formalized in the following Theorem.
Theorem 2.5.
Assume that for at least one $i$, $\psi_{i}\neq 0_{k}$ but at least one element of $\psi_{i}$ is null, i.e. $\psi_{ih}=0$. Let $P\neq I$ be any rotation matrix such that $P_{hl}\neq\{0,\pm 1\}$ (i.e. $P$ is not a permutation matrix), let $\Psi$ be as in Definition 2.1, $\Phi$ be the $n\times k$ matrix with the vectors $\varphi_{i}$ organized as its rows, $\tilde{\Phi}$ the $n\times k$ matrix that similarly stacks the $\tilde{\varphi}_{i}$, such that $\Phi=\Psi\odot\tilde{\Phi}$, where $\odot$ denotes the Hadamard matrix multiplication. Then, $\Phi\Gamma^{\top}\neq[\Psi\odot(\tilde{\Phi}P)]P^{\top}\Gamma^{\top}$.
@@
The theorem is proved by contradiction. Assume $[\Psi\odot\tilde{\Phi}]=[\Psi\odot(\tilde{\Phi}P)]P^{\top}.$ If we multiply both sides of this equality by the matrix $P$, we get $[\Psi\odot\tilde{\Phi}]P=[\Psi\odot(\tilde{\Phi}P)].$ We can write the generic element in row $i$, column $h$ in the left hand side of equation above as
$\{[\Psi\odot\tilde{\Phi}]P\}_{ih}=[\psi_{i}\odot\tilde{\varphi}_{i}]^{\top}P_{h}=\sum_{l=1}^{k}\psi_{il}\tilde{\varphi}_{il}P_{lh},$
and each element of the right hand side as meet $[\Psi\odot(\tilde{\Phi}P)]_{ih}=\psi_{ih}\sum_{l=1}^{k}\tilde{\varphi}_{il}P_{hl}.$ In particular, since for at least one $i$, there exist a $\psi_{ih}=0$ and $\psi_{i}\neq 0_{k}$ the equality holds true only if $P_{h}$ is a suitable combination of 0 or 1 but this is impossible as $P$ is not a permutation matrix. ∎
Note that the strongest condition in Theorem 2.5 is that there exist at least one observation $i$ with $\psi_{i}\neq 0_{k}$ and at least one element of $\psi_{i}$ is null. This means, for example that if there is one group with no specific factors and other groups with specific factor, the identification of $\Lambda$ is not ensured. In practice, one can check a posteriori whether this condition is satisfied or not before trying to interpret the results.
@@
Clearly, the shared variability induced by $\Lambda$ remains susceptible to well-known identifiability issues. However, these challenges can be effectively addressed using state-of-the-art methods, such as those proposed by [4] and [7].
2.3 Posterior computation
In this section, we present a Markov Chain Monte Carlo scheme to sample from the posterior distribution of the parameters of APAFA. We first rewrite the model as
$y_{ij}=\sum_{h=1}^{d}{\lambda}_{jh}\eta_{ih}\tau_{h}+\sum_{h=1}^{k}\tilde{\varphi}_{ih}\psi_{ih}\tau^{\varphi}_{h}+\epsilon_{ij},$ (9)
with $\varphi_{ih}=\psi_{ih}\tilde{\varphi}_{ih}$ with $\tilde{\varphi}_{ih}>0$. The parameters are then updated sequentially iterating the following steps.
  1. 1.
    Update the block of latent factors $[\eta_{i},\tilde{\varphi}_{i}]$ for each $1\leq i\leq n.$
    • Sample from $[\eta_{i},\tilde{\varphi}_{i}]\sim N_{d+k}\left(\hat{\mu}_{\eta,\varphi},\hat{\Sigma}_{\eta,\varphi}\right),$ with
      $\hat{\Sigma}_{\eta,\varphi}=\left(I_{d+k}+([\Lambda\ ,\Gamma I_{k}(\psi_{i}\odot\tau^{\varphi})])^{\top}\Sigma^{-1}[\Lambda,\Gamma I_{k}(\psi_{i}\odot\tau^{\varphi})]\right)^{-1},$
      $\hat{\mu}_{\eta,\varphi}=\hat{\Sigma}_{\eta,\varphi}([\Lambda,\Gamma I_{k}(\psi_{i}\odot\tau^{\varphi})])^{\top}\Sigma^{-1}y_{i},$
      where $a\odot b$ indicates the term-by-term product of vectors $a$, $b$, $I_{k}(\cdot)$ the identity matrix, and $\varphi_{i}=\tilde{\varphi}_{i}\odot\psi_{i}$.
    • Standardize the factors computing $\bar{\eta}_{h}=n^{-1}\sum_{i=1}^{n}\eta_{ih}$ $V^{\eta}_{h}=n^{-1}\sum_{i=1}^{n}(\eta_{ih}-\bar{\eta_{h}})^{2}$ for each $h=1,\ldots,d$ , and $\bar{\tilde{\varphi}}_{h}=n^{-1}\sum_{i=1}^{n}\tilde{\varphi}_{ih}$, $V^{\varphi}_{h}=n^{-1}\sum_{i=1}^{n}(\tilde{\varphi_{ih}}-\bar{\tilde{\varphi_{h}}})^{2},$ for $h=1,\ldots,k$. Then set $\eta_{h}=\eta_{h}/\sqrt{V^{\eta}_{h}}$, $\Lambda_{h}=\Lambda_{h}\sqrt{V^{\eta}_{h}}$, $\tilde{\varphi}_{h}=\tilde{\varphi}_{h}/\sqrt{V^{\varphi}_{h}}$, $\varphi_{h}=\varphi_{h}/\sqrt{V^{\varphi}_{h}}$, $\Gamma_{h}=\Gamma_{h}\sqrt{V^{\varphi}_{h}}$.
  2. 2.
    Update $\Sigma$ by sampling the diagonal elements for $j=1,\ldots,p$,
    $\sigma_{j}^{2}\sim\mathrm{IGa}\left\{a_{\sigma}+\frac{n}{2},b_{\sigma}+\frac{1}{2}\sum_{i=1}^{n}\left(y_{ij}-\lambda_{j}^{\top}\eta_{i}-\gamma_{j}^{\top}\varphi_{i}\right)^{2}\right\}.$
  3. 3.
    Update $\beta_{h}=(\beta_{h1},\ldots,\beta_{hs})^{\top}$ for $h=1,\ldots,k.$ as follows
    • for $h=1,\ldots,k$, sample $n$ independent Pólya-Gamma distributed random variables [18] $\omega_{ih}\sim\mathrm{PG}\left(1,x_{i}^{\top}\beta_{h}\right)$
    • Let $D_{h}$ denote the $n\times n$ diagonal matrix with entries $d_{ih}(i=1,\ldots,n)$ and let $B=\sigma_{\beta}^{2}I_{S}$ be a $S\times S$ diagonal matrix. For each $h=1,\ldots,H$, update $\beta_{h}$ sampling
      $\beta_{h}\sim N_{S}\left\{\left(x^{\mathrm{T}}D_{h}x+B^{-1}\right)^{-1}\left(x^{\mathrm{T}}\kappa_{h}\right),\left(x^{\mathrm{T}}D_{h}x+B^{-1}\right)^{-1}\right\},$
      where $\kappa_{h}$ is a $p$-dimensional vector with the $j$-th entry equal to $\psi_{ih}-0.5$.
  4. 4.
    Update the variance of the loadings
    $\Sigma^{\lambda}_{hh}\sim\text{IGa}(a^{\lambda}+n/2,b^{\lambda}+1/2\sum_{j=1}^{p}\lambda_{jh}^{2}),$
    $\Sigma^{\gamma}_{hh}=\text{IGa}(a^{\gamma}+n/2,b^{\gamma}+1/2\sum_{j=1}^{p}\gamma_{jh}^{2}).$
  5. 5.
    Update the loadings $\lambda_{j}$ and $\gamma_{j}$ (the rows of loading matrix) by sampling from the independent full conditional : define $\Sigma^{*}$ as the $(d+k)\times(d+k)$ covariance matrix that collects the diagonal elements of $\Sigma^{\lambda}$ and $\Sigma^{\gamma}$ and $y_{j}$ the $j$-th columns of $y$
    $[\lambda_{j},\gamma_{j}]\sim N_{d+k}\{(\Sigma^{-1*}+[\eta_{j}\odot\tau,\varphi_{j}\odot\tau^{\varphi}]^{\top}\sigma_{j}^{-2}[\eta_{j}\odot\tau,\varphi_{j}\odot\tau^{\varphi}])^{-1}[\eta_{j}\odot\tau,\varphi_{j}\odot\tau^{\varphi}]^{\top}\sigma_{j}^{-2}y_{j},\\ \left(\Sigma^{-1*}+[\eta_{j}\odot\tau,\varphi_{j}\odot\tau^{\varphi}]^{\top}\sigma_{j}^{-2}[\eta_{j}\odot\tau,\varphi_{j}\odot\tau^{\varphi}]\right)^{-1}\}.$
  6. 6.
    Update the cumulative shrinkage parameters for the number of shared factors $\tau$. Define indicators $z^{\eta}_{h}$ with prior $pr(z^{\eta}_{h}=l)=w^{\eta}_{l}$ for $l=1,\ldots,d$. Sample from
    $\operatorname{pr}\left(z^{\eta}_{h}=l\right)\propto\begin{cases}w^{\eta}_{l}\prod_{i=1}^{n}\prod_{j=1}^{p}N\left(y_{ij};\mu_{ijh}^{(0)},\sigma_{j}^{2}\right)&\text{ for }\quad l=1,\ldots,h\\ w^{\eta}_{l}\prod_{i=1}^{n}\prod_{j=1}^{p}N\left(y_{ij};\mu_{ijh}^{(1)},\sigma_{j}^{2}\right)&\text{ for }\quad l=h+1,\ldots,d,\end{cases}$
    where $\mu_{ijh}^{(z)}=\sum_{l\neq h}^{d}\sqrt{\tau_{l}}\lambda_{jl}\eta_{il}+z\lambda_{jh}\eta_{ih}+\sum_{l=1}^{k}\sqrt{\tau^{\varphi}_{l}}\sqrt{\psi_{il}}\gamma_{jl}\varphi_{il}$. Set $\tau_{h}=1$ if $z^{\eta}_{h}>h$, else $\tau_{h}=0$.
  7. 7.
    Update the cumulative shrinkage parameters for the number of specific factors $\tau^{\varphi}$. Define indicators $z^{\varphi}_{h}$ with prior $pr(z^{\varphi}_{h}=l)=w^{\varphi}_{l}$ for $l=1,\ldots,k$. Sample from
    $\operatorname{pr}\left(z^{\varphi}_{h}=l\right)\propto\begin{cases}w^{\varphi}_{l}\prod_{i=1}^{n}\prod_{j=1}^{p}N\left(y_{ij};\mu_{ijh}^{(0)},\sigma_{j}^{2}\right)&\text{ for }\quad l=1,\ldots,h\\ w^{\varphi}_{l}\prod_{i=1}^{n}\prod_{j=1}^{p}N\left(y_{ij};\mu_{ijh}^{(1)},\sigma_{j}^{2}\right)&\text{ for }\quad l=h+1,\ldots,k,\end{cases}$
    where $\mu_{ijh}^{(z)}=\sum_{l=1}^{d}\sqrt{\tau_{l}}\lambda_{jl}\eta_{il}+\sum_{l\neq h}^{k}\sqrt{\tau^{\varphi}_{l}}\sqrt{\psi_{il}}\gamma_{jl}\varphi_{il}+z\sqrt{\psi_{ih}}\gamma_{jh}\varphi_{ih}$. Set $\tau^{\varphi}_{h}=1$ if $z^{\varphi}_{h}>h$, else $\tau^{\varphi}_{h}=0$.
  8. 8.
    For $l=1,\ldots,d-1$, sample $v_{l}^{\eta}$ from
    $\left(v_{l}^{\eta}\mid-\right)\sim\operatorname{Be}\left\{1+\sum_{h=1}^{d}\mathbf{1}\left(z_{h}^{\eta}=l\right),\alpha^{\eta}+\mathbf{1}\left(z^{\eta}_{h}>l\right)\right\},$
    set $v^{\eta}_{d}=1$ and update $w^{\eta}_{l}=v^{\eta}_{l}\prod_{m=1}^{l-1}\left(1-v^{\eta}_{m}\right)$, for $l=1,\ldots,d$. Proceed analogously for $v^{\varphi}_{l}$ and $w^{\varphi}_{l}$.
  9. 9.
    Update $\psi_{ih}$, for $i=1,\ldots,n$ and $h=1,\ldots,k$ by sampling from the full conditional distributions
    $\operatorname{pr}\left(\psi_{ih}=u\right)\propto\begin{cases}\left\{1-\operatorname{logit}^{-1}\left(x_{i}^{\top}\beta_{h}\right)\right\}\prod_{j=1}^{p}N\left(y_{ij};\mu_{ijh}^{(u)},\sigma_{j}^{2}\right)&\text{ for }u=0\\ \operatorname{logit}^{-1}\left(x_{i}^{\top}\beta_{h}\right)\prod_{j=1}^{p}N\left(y_{ij};\mu_{ijh}^{(u)},\sigma_{j}^{2}\right)&\text{ for }u=1\end{cases}$
[…]
  10. 10.
    Reorder the columns of $\Gamma$ and $\Lambda$ and factors $\eta,$ $\phi,$ and $\psi$ according to active and non active components of $\tau$ and $\tau^{\varphi}$. Move to shared components factors s.t. $\psi_{h}=1_{n}$.

### v2

@@ 2.2 Model properties
Identifiability of factor models is an overarching issue. While these models are appealing in revealing possible interpretable relationships between a small set of latent factors and multivariate observations—even with complex dependency structures—this interpretability comes at a cost: factor analytic models lack identifiability in fundamental ways. For instance, loadings are not identifiable due to rotation, sign and permutation invariance. Rotational invariance refers to the transformation of loadings and factors by an orthogonal matrix, sign switching arises because one can simultaneously change the signs of certain elements without affecting the overall structure, while permutation invariance occurs because there is no intrinsic ordering in the factor loading matrix columns. See Papastamoulis & Ntzoufras (2022) Xu et al. (2023) Frühwirth-Schnatter et al. (2024) for a comprehensive review.
@@
Identification issues are exacerbated in the MSFA framework of De Vito et al. (2021) where it may be difficult to disentangle the signal between shared and study-specific factors, a problem known as information switching. For instance, Chandra et al. (2024) shows how a linear combination of study-specific factors can instead be represented and reincorporated as part of the shared component. Building on these concepts, we provide conditions to identify the shared from the specific factors in our setup. Throughout, let $\mathcal{O}_{k}$ denote the space of $k\times k$ orthogonal matrices. We formalize the notion of information switching through the following definition:
Definition 1.
@@
Let $n$ be the number of units, and $S_{n}$ the number of distinct groups in model (7), i.e. $S_{n}=\left|\cup_{i=1}^{n}\Omega_{i}\right|$. Denote with $\Psi$ the $n\times k$ matrix that stacks in distinct rows all $\psi_{i}=(\psi_{i1},\ldots,\psi_{ik})$ and with $\Psi_{h}$ its generic column with, $\Psi_{h}\neq 1_{n}$ for all $h=1,\dots,k$. Let $\Omega_{s}^{*}$ and $\psi_{s}^{*}$ ($s=1,\dots,S_{n}$) be the distinct values of $\Omega_{i}$ and $\psi_{i}$, respectively. Similarly, let $W_{s}^{*}=\Omega_{s}^{*}-\Lambda\Lambda^{\top}-\Sigma=\Gamma\text{diag}\{\psi_{s}^{*}\}\Gamma^{\top}$. The model suffers from information switching if there exist $\tilde{\Gamma}\neq\Gamma$ and $\tilde{\Psi}\neq\Psi$ such that $W_{s}^{*}=\tilde{\Gamma}{\mbox{diag}\{\tilde{\psi}_{s}^{*}\}}\tilde{\Gamma}$ for all $s$, with $\tilde{\Psi}_{h}=1_{n}$ for at least one $h$.
@@
It is clear that if $\tilde{\Psi}_{h}=1_{n}$, the $h$-th factor could no longer be interpreted as specific and should be moved to the shared part. Consistently, $\Lambda\Lambda^{\top}$ would represent only a fraction of the shared variance. Note that in Definition 1, $S_{n}$ does not refer to the number of distinct labels or studies $S$ fixed a priori and embedded in the categorical variables $x_{i}$, but rather it is equal to the number of distinct configurations of the rows of the matrix $\Psi$. The following theorem ensures the identification in terms of information switching.
Theorem 1.
@@
For the model defined in (7), if $\Psi_{h}\neq 1_{n}$ for all $h\in\{1,\ldots,k\}$ and $\Gamma$ is of full column rank $k$ with $k<p(p+1)/2$, then the model is resistant to information switching.
@@
Let $\Psi^{*}$ be the $S_{n}\times k$ matrix containing the distinct $\psi^{*}_{s}$. Since each $\psi_{s}^{*}$ is a vector of zeroes and ones, we can see that $W_{s}^{*}=\Gamma\text{diag}\{\psi_{s}^{*}\}\Gamma^{\top}=\Gamma\text{diag}\{\psi_{s}^{*}\}\text{diag}\{\psi_{s}^{*}\}\Gamma^{\top}.$ As in any factor analytic decompositions, if $P_{s}\in\mathcal{O}_{k}$ we can write $W^{*}_{s}=\Gamma\text{diag}\{\psi_{s}^{*}\}P_{s}P_{s}^{\top}\text{diag}\{\psi_{s}^{*}\}\Gamma^{\top}.$ Consistently, each element $w^{*(jl)}_{s}$ of $W_{s}$, for $j,l=1,\ldots,p$ can be written as
$w^{*(jl)}_{s}=\sum_{h=1}^{k}\gamma_{jh}\gamma_{lh}{\bar{\psi}_{hs}\bar{\psi}_{hs}^{\top}},$
where $\bar{\psi}_{hs}$ is the $h$-th column of the matrix product $\text{diag}\{\psi_{s}^{*}\}P_{s}.$ Since $P_{s}\in\mathcal{O}_{k}$, the product $\bar{\psi}_{hs}\bar{\psi}_{hs}^{\top}$ is either 0 or 1 by construction. Thus, we know that there exist binary variables defined as $\tilde{\psi}^{*}_{hs}=\bar{\psi}_{hs}\bar{\psi}_{hs}^{\top}$ for $h=1,\dots k$ leading to the same $W_{s}^{*}$ for each $s=1,\dots,S_{n}$. Hence, it is sufficient to check whether these $\tilde{\psi}_{s}^{*}$ are actually different from the original $\psi_{s}^{*}$ for each $s$. To this end, for the pair $(jl)$ consider
$\sum_{h=1}^{k}\gamma_{jh}\gamma_{hl}\tilde{\psi}_{hs}^{*}=w^{*(jl)}_{s}$
as a linear equation in $\tilde{\psi}_{hs}^{*}$. Since we have $S_{n}p(p+1)/2$ linear equations of this type, we can frame them into a linear system $G\mathbf{x}=\mathbf{w}$ defying $G$ as the $S_{n}p(p+1)/2\times k$ matrix of known coefficients, $\mathbf{x}$ as the vector of variables, here associated to the $\tilde{\psi}_{hs}^{*}$, and $\mathbf{w}$ as the vector stacking the half-vectorization of $W_{s}^{*}$ (containing its upper triangular entries) for $s=1,\dots,S_{n}$. Specifically, the full design matrix $G=I_{S_{n}}\otimes G_{s}$, where $G_{s}$ is the $p(p+1)/2\times k$ matrix whose rows correspond to index pairs $(j,l)$ with $1\leq j\leq l\leq p$, and whose entries are given by
$G_{(j,l)h}=\gamma_{jh}\gamma_{lh},\quad\text{for }h=1,\dots,k.$
Since $\text{rank}(G_{s})=k$, then $\text{rank}(G)=S_{n}k$, and $\text{rank}(G)=\text{rank}(G|\mathbf{w})$ by construction. Thus, since $k<p(p+1)/2$ by assumption, the solution is unique by the Rouché–Capelli Theorem. The uniqueness of the solution for the elements $\psi_{s}^{*}$ ensures that it is impossible to find a representation where for one $h$, $\tilde{\Psi}_{h}$ is a vector of one.
The proof of Theorem 1 is based on proving existence of unique vectors $\psi_{s}^{*}$ for the specific factors. This guarantees that no alternative representation exists in which, for some $h$, the corresponding $\tilde{\Psi}_{h}$ is a vector of ones. Notably, it also establishes the uniqueness of the sparsity patterns encoded by the $\Psi$ matrix. That is, any alternative representation inducing the same covariance structure must preserve the subject-specific activation patterns of the specific components. The following corollary formalizes this consequence.
Corollary 1.
Under the conditions of Theorem 1, the sparsity indicators $\psi^{*}_{hs}$ are uniquely identified up to a permutation of the specific factors.
Based on the identifiability findings of Theorem 1, the practical conditions that ensures resistance to information switching in APAFA are discussed in Lemma 1 and 2.
Lemma 1.
@@
A key advantage of our model is that the requirement of $k<p(p+1)/2$ is minimal as it is customary in factor analysis to have latent factors of a considerably lower dimension than the observed data ($k<p$). Notably, Chandra et al. (2024), in different settings, necessitate the dimension of the specific part to be smaller than that of the shared part implicitly considering the study-specific part as a noise or disturbance parameter. Although we are in different settings, where the number and composition of the groups is not pre-specified, we do not restrict $k$ to be also smaller than $d$. This is crucial in many contexts. For instance, in cancer studies, it is possible to study very different types of cancer simultaneously, identifying a range of specific factors that describe the specific characteristics of each type of cancer, while also uncovering a few common factors that are shared across cancers, regardless of the organ or tissue affected possibly pertaining to fundamental mechanisms of oncogenesis.
Under stronger assumptions, detailed below, we show that the specific factors are not only distinguishable from the shared ones, but also fully identifiable up to a permutation of their columns. We start providing the following definition.
Definition 2 (Non-replicable Sparsity Pattern Condition).
All columns of $\Psi^{*}$, where $\Psi^{*}$ is the $S_{n}\times k$ matrix stacking all the distinct $\psi^{*}_{s}$, are different.
We now provide a result that, under the non-replicable sparsity pattern condition, ensures full identifiability of the specific factor, up to permutation.
Theorem 2.
Let $\Gamma\in\mathbb{R}^{p\times k}$ be a real matrix with full column rank $k.$ If $P\in\mathcal{O}_{k}$ and $\Gamma^{\prime}=\Gamma P$ is a rotation of the specific factors, under the Non-replicable Sparsity Pattern Condition (Definition 2), then
$\Lambda\Lambda^{\top}+\Gamma\Gamma^{\top}+\Sigma=\Lambda\Lambda^{\top}+\Gamma^{\prime}\Gamma^{\prime\top}+\Sigma,$
and, for each $s=1,\dots,S_{n}$
$\Lambda\Lambda^{\top}+\Gamma\Psi_{s}^{*}\Gamma^{\top}+\Sigma=\Lambda\Lambda^{\top}+\Gamma^{\prime}\mbox{diag}(\psi_{s}^{*})\Gamma^{\prime\top}+\Sigma,$
if and only if $P$ is a permutation matrix.
@@
We begin by fixing the ordering of the columns of $\Gamma$. Under this fixed ordering, we will show that the only admissible orthogonal matrices $P\in\mathcal{O}_{k}$ are diagonal with entries ±$\pm 1$. Notably, applying any permutation to the columns of $\Gamma$ initially and repeating the argument would yield the same conclusion, i.e. the only admissible $P$ relative to the permuted ordering are diagonal with entries ±$\pm 1$. Consequently, the general solution requires $P$ to be a signed permutation matrix.
Let $A_{s}$ be the active support of group $s$, namely
$A_{s}:=\{h\in\{1,\dots,k\}\,:\,\psi^{*}_{hs}=1\},$
and define $\Gamma_{s}\in\mathbb{R}^{p\times k_{s}}$ to be the submatrix of $\Gamma$ consisting of the columns indexed by $A_{s}$. Notably these are identifiable thanks to Corollary 1.
The proof proceeds considering that $P_{s}\in\mathcal{O}_{k_{s}}$, leaves $\Gamma^{\prime}_{s}\Gamma^{\prime}_{s}{}^{\top}$ invariant and hence that the marginal distribution of group $s$ is also invariant under such rotations. Given these $P_{s}\in\mathcal{O}_{k_{s}}$ for each $s=1,\dots,S_{n}$, we investigate whether a “global” orthogonal matrix $P\in\mathcal{O}_{k}$ can be constructed by assembling the blocks $P_{s}$. This requires consistency across overlapping blocks, ensuring that each $P_{s}$ remains orthogonal, for example satisfying $\sum_{l\in A_{s}}\|P_{s,hl}\|=1$ for all $h$ and $s$. We will argue in the following that, under the Non-replicable Sparsity Pattern Condition of Definition 2, this is possible only if $P$ is a diagonal matrix with entries ±$\pm 1$.
For any $h=1,\dots,k$ there exist at least one group loading factor $h$, i.e. $\psi^{*}_{sh}=1$ for at least one $s$. Then we have two possibilities: $i)$ $\psi_{s^{\prime}h}^{*}=0$ for all other $s^{\prime}\neq s$, or $ii)$ $\psi_{s^{\prime}h}^{*}=1$ for at least another $s^{\prime}\neq s$.
In the first case $i)$, $h$ is factor belonging exclusively to group $s$. Notably, by Definition 2 it is the only exclusive factor. If this is the only factor of group $s$, then the rotation matrix $P_{s}$ reduces to a scalar. Since $P$ is orthogonal, $\|P_{h}\|^{2}=1$ implies $P_{hh}=\pm 1$ and all other entries in the $h$th row/column of $P$ must vanish. If $h$ is not the only factor of group $s$, then $A_{s}$ contains at least another $h^{\prime}$ that is shared among $s$ and at least another $s^{\prime}$. In this case, $h^{\prime}$ is in case $ii)$. Consider then case $ii)$. In order to be orthogonal, matrix $P$ needs to satisfy
 $\>1=\|P_{h}\|^{2}=\sum_{l=1}^{k}P_{lh}^{2}=\sum_{l\in A_{s^{\prime}}\cap A_{s}}P_{lh}^{2}+\sum_{l\in A_{s^{\prime}}\setminus A_{s}}P_{lh}^{2}+\sum_{l\in A_{s}\setminus A_{s^{\prime}}}P_{lh}^{2}+\sum_{l\notin A_{s}\cup A_{s^{\prime}}}P_{lh}^{2}.$ (8)
Additionally, since $h\in A_{s}$, then
 $\>1=\|P_{{s},h}\|^{2}=\sum_{l\in A_{s}\cap A_{s^{\prime}}}P_{lh}^{2}+\sum_{l\in A_{s}\setminus A_{s^{\prime}}}P_{lh}^{2}$ (9)
and similarly, since $h\in A_{s^{\prime}}$,
 $\>1=\|P_{{s^{\prime}},h}\|^{2}=\sum_{l\in A_{s^{\prime}}\cap A_{s}}P_{lh}^{2}+\sum_{l\in A_{s^{\prime}}\setminus A_{s}}P_{lh}^{2}.$ (10)
Equations (8), (9), and (10) enforce that all $P$, $P_{s}$, and $P_{s^{\prime}}$ are orthogonal matrices. These set of equations take the form: $a+b+c+d=1,\>a+b=1,\>a+c=1$ with $a,b,c,d\geq 0$. This system admits only the solution $a=1$ and $b=c=d=0$. This observation directly constrains the entries of the $h$-th column of $P$ pertaining to the last three summations of (8) to be zero. Consequently, the only actual rotations, i.e. those induced by non-diagonal $P_{s}$, are those corresponding to factors shared between groups $s$ and $s^{\prime}$, i.e. the $h$-th column of $P$ must vanish outside $A_{s}\cap A_{s^{\prime}}$.
Assumption 2 induces further constraints on $P$. Specifically, for any pair $(h,h^{\prime})$ jointly in groups $s$ and $s^{\prime}$, there exists a third group $s^{\prime\prime}$ such that either $h\in A_{s^{\prime\prime}}$ but $h^{\prime}\notin A_{s^{\prime\prime}}$, or $h^{\prime}\in A_{s^{\prime\prime}}$ but $h\notin A_{s^{\prime\prime}}$. Applying the same norm argument to $s$ and $s^{\prime\prime}$ (or equivalently to $s^{\prime}$ and $s^{\prime\prime}$) forces the $h$th column of $P$ to vanish outside $A_{s}\cap A_{s^{\prime\prime}}$. Crucially, since $A_{s}\cap A_{s^{\prime\prime}}$ excludes $h^{\prime}$ (or $h$), we conclude $P_{hh^{\prime}}=P_{h^{\prime}h}=0$ and $P_{hh}=P_{h^{\prime}h^{\prime}}=\pm 1$. The argument is valid for all $h$, leading to $P$ being a a diagonal matrix with entries ±$\pm 1$. ∎
@@
Clearly, the shared variability induced by $\Lambda$ remains susceptible to well-known identifiability issues. However, these challenges can be effectively addressed using state-of-the-art methods, such as those proposed by Poworoznek et al. (2021) and Papastamoulis & Ntzoufras (2022).